# Simulating the Galactic population of axion clouds around stellar-origin black holes: Gravitational wave signals in the 10 - 100 kHz band

### v2

@@ Abstract
Ultralight scalar fields can experience runaway ‘superradiant’ amplification near spinning black holes, resulting in a macroscopic ‘axion cloud’ which slowly dissipates via continuous monochromatic gravitational waves. For a particular range of boson masses, $\mathcal{O}(10^{-11}$$10^{-10})$ eV, an axion cloud will radiate in the 10101010100100100100 kHz band of the Levitated Sensor Detector (LSD). Using fiducial models of the mass, spin, and age distributions of stellar-origin black holes, we simulate the present-day Milky Way population of these hypothetical objects. As a first step towards assessing the LSD’s sensitivity to the resultant ensemble of GW signals, we compute the corresponding signal-to-noise ratios which build up over a nominal integration time of $10^{7}$ s, assuming the projected sensitivity of the $1$-m LSD prototype currently under construction, as well as for future $10$-m and $100$-m concepts. For a $100$-m cryogenic instrument, hundreds of resolvable signals could be expected if the boson mass $\mu$ is around $3\times 10^{-11}$ eV, and this number diminishes with increasing $\mu$ up to $\approx 5.5\times 10^{-11}$ eV. The much larger population of unresolved sources will produce a confusion foreground which could be detectable by a $10$-m instrument if $\mu\in(3-4.5)\times 10^{-11}$ eV, or by a $100$-m instrument if $\mu\in(3-6)\times 10^{-11}$ eV.
@@ I Introduction
This scenario can be realized with physics beyond-the-Standard-Model (BSM). For example, a large number of ultralight fields may occur as a result of the compactification of extra dimensions [17]. One of these may be the QCD axion – the pseudoscalar boson proposed to solve the strong-CP problem [18][19][20]. The axion is a Goldstone boson of a spontaneously-broken global symmetry which acquires a small mass through non-perturbative effects. Its mass, $\mu$, is determined by the energy scale $f_{a}$ associated with the broken symmetry [17],
@@
The $1-4\ M_{\odot}$ range of BH masses is gradually being populated by microlensing candidates, X-ray binary candidates, and GW events such as GW190814 [7] and GW230529 [8]. Since the mass distribution for these objects is still unknown, we will limit our attention to stellar-origin BH’s with masses between $5$ and $20\ M_{\odot}$, typical of BH’s in X-ray binary systems.
@@ II Superradiant bound-states
which is the radial Schrödinger equation for a non-relativistic particle in a Coulomb potential $V(r)=\alpha/r$ – hence the moniker ‘gravitational atom’. The $\alpha^{2}/(2n^{2})$ term in Eq. II is precisely the ‘hydrogen atom’ solution to Eq. 10. The complete small-$\alpha$ solutions to the full Klein-Gordon equation have been computed order-by-order using the method of matched asymptotic expansions [16],
@@ III Gravitational waves from axion clouds
At a particle-physics level, GW production by axion clouds can be understood in terms of two processes: annihilation of two bosons to a single graviton (with the BH absorbing the recoil momentum), and downward-transitions between bound-states [31]. However, just as superradiance is a purely classical kinematic effect, the GW emission can also be understood classically in terms of the cloud’s time-dependent quadrupole moment. That being said, the GW signals considered in this work correspond to the annihilation channel.
@@
where $F_{+}(t)$ and $F_{\crossproduct}(t)$ are the detector’s angular pattern functions. The amplitudes $a_{+/\crossproduct}$ are expanded in terms of spheroidal harmonics with spin-weight $s=-2$,
@@
Comparing Eqs. 21 and 22, we see that $\mathcal{A}_{\tilde{l}\tilde{m}}\propto\sqrt{d\tilde{E}/dt}$, allowing us to express $h_{0}^{(\tilde{l})}$ directly in terms of $d\tilde{E}/dt$. Restricting ourselves to the dominant mode $\tilde{m}=\tilde{l}=2l$, we obtain a closed-form solution for the characteristic amplitude, which we use without abandon to compute the GW amplitudes of the axion clouds resulting from our simulations (we will drop the superscript $(2l)$ henceforth),
@@ V Simulation procedure
Those BH’s with an extant cloud are assigned a location in the Milky Way (Eqs. 30 and 31). Earth is assigned to an arbitrary, but fixed, point on the circle of radius $8.3$ kpc in the Galactic midplane. For a cloud located at distance $d$, we check the inequality $ct_{r}>d$ to determine if there has been enough time for GW’s to propagate to Earth since the cloud formed. Those clouds for which $d>ct_{r}$ are presently unobservable, and we retain only those clouds for which $ct_{r}>d$. We summarize this section with the following flowchart:
@@ VI.2 Resolvable signals
where $\sqrt{S_{n}(f)}$ is the one-sided amplitude spectral density (ASD) of the detector noise (the ‘sensitivity curve’) evaluated at the GW frequency, and the trapping frequency of the levitated sensor is constant during the entire observation time. Although the LSD is an Earth-bound detector for which the observed signal is modulated by the Earth’s daily (diurnal) rotation and orbital motion, we, instead, compute the SNR for the idealized case of a detector freely orbiting the Milky Way at the same radius as the Solar System (i.e. not attached to a planet or star system). This scenario isolates the intrinsic sensitivity of the LSD to continuous-wave signals from incidental factors, such as the Earth’s rotational and orbital periods.
@@
The ‘loudness’ of a signal is determined primarily by the source distance. The distance, in turn, is a random variable determined by the randomly-assigned position vector (Eqs. 30 and 31) of the source. Thus, for a given set of extant clouds, the number of individually-resolved sources $N_{\text{res}}$ will vary each time we re-assign their position vectors. We estimate the mean & standard deviation of $N_{\text{res}}$ for a given population of extant clouds by laying them down in the Galaxy $N_{\text{reshuffle}}=100$ times and counting how many are resolvable in each ‘re-shuffling’. The mean & standard deviation are then computed as
@@ VI.3 Unresolved signals
A preliminary method for estimating the LSD’s sensitivity to the confusion foreground is to treat each pair $(f_{c},\ h_{\text{eff}})$ as if they were the frequency and amplitude of a hypothetical monochromatic signal whose corresponding effective SNR $\rho_{\text{eff}}$, computed via Eq. 33, is then compared to a threshold $\rho_{\text{t}}$. We continue to require $\rho_{\text{t}}=10$. The numerator and denominator of Eq. 33 $\big{(}h_{\text{eff}}\sqrt{T_{\text{coh}}}\ \text{and}\ \sqrt{S_{n}(f_{c})},\ \text{respectively}\big{)}$ are shown separately in Fig. 10, and their ratio (the SNR) is shown in Figs. 11, 12, and 13 for the $1-$, $10-$, and $100-$m instruments, respectively.
@@ VII Conclusion
Finally, we note the following limitations of this work, as well as directions for future work: First, since isolated BH’s have no EM counterpart, we do not know, ahead of time, the direction to these GW sources. Targeted & directed searches for axion clouds will, therefore, not be possible for isolated BH’s, and we must resort to blind searches. A full template search will require us to pixelate the sky, and it is an open question what is the smallest angular size $\Delta\theta$ for which the number of templates in a blind all-sky search is not prohibitively large: We do not want the time required for data analysis to become larger than the four-month observation period. Transverse proper motions will also need to be accounted for if, over the observation time, a source moves out of the pixel on the sky it was in initially.
Diurnal and annual modulations of the GW frequency & amplitude must also be included in our detection scheme. The angular dependence of the detector sensitivity, encoded in the detector’s pattern functions, produces a daily modulation of the signal amplitude. The rotational and orbital motions of the Earth produce time-varying Doppler shifts. The Doppler modulation can be corrected, but doing so requires knowledge of the source position to a precision determined by the observation time and GW frequency.
In this work, we have used the SNR as a baseline detection statistic. We have not yet determined what are the most appropriate detection statistics & criteria for the monochromatic signals from individual clouds. Moreover, our treatment of the confusion foreground has not accounted for the intrinsic anisotropy of the signal: The axion clouds will be distributed throughout the disks and bulge of the Milky Way, so the strength of the foreground will vary over the sky in a complicated way. Additionally, searches for stochastic signals typically involve an ‘excess-power’ method, as well as cross-correlation between multiple detectors. Plans to build a second $1$-m instrument at UC Davis (in addition to the Northwestern detector) are in development, so while a single $1$-m detector might not have the requisite sensitivity, the prospects for a two- or multi-detector scheme are an exciting avenue of future study.
@@ Appendix A Superradiant bound-states
Strictly speaking, the radial and angular eigenvalues, $\xi$ and $\Lambda$, must be found simultaneously. Leaver’s method can also be applied to 47 [50], resulting in a continued-fraction condition analogous to 64. We then have two equations for the two unknowns.
Conveniently, we can reduce the root-finding problem to merely solving 64 by using the $Mathematica$ function ‘SpheroidalEigenvalue’. With the change of variable $z=\cos\theta$, and in terms of the following quantities,
@@
For our axion cloud simulations, we have needed to compute $\xi$ for bound-states up to, and including, $n=9$. As an example, we have plotted the real & imaginary parts of the $n=8$ bound-state in Figs. 14, 15 and 16; Figs. 14 and 15 are analogous to Figs. 1 and 2.

### v4

# Simulating the Galactic population of axion clouds around stellar-origin black holes: Gravitational wave signals in the 10 – 100 kHz band

@@ Abstract
Ultralight scalar fields can experience runaway ‘superradiant’ amplification near spinning black holes, resulting in a macroscopic ‘axion cloud’ which slowly dissipates via continuous monochromatic gravitational waves. For a particular range of boson masses, $\mathcal{O}(10^{-11}$$10^{-10})$ eV, an axion cloud will radiate in the 10101010100100100100 kHz band of the Levitated Sensor Detector (LSD). Using fiducial models of the mass, spin, and age distributions of stellar-origin black holes, we simulate the present-day Milky Way population of these hypothetical objects. As a first step towards assessing the LSD’s sensitivity to the resultant ensemble of gravitational wave signals, we compute the corresponding signal-to-noise ratios which build up over a nominal integration time of $10^{7}$ s, assuming the projected sensitivity of the $1$-m LSD prototype currently under construction, as well as for future $10$-m and $100$-m concepts. For a $100$-m cryogenic instrument, hundreds of resolvable signals could be expected if the boson mass $\mu$ is around $3\times 10^{-11}$ eV, and this number diminishes with increasing $\mu$ up to $\approx 5.5\times 10^{-11}$ eV. The much larger population of unresolved sources will produce a confusion foreground which could be detectable by a $10$-m instrument if $\mu\in(3-4.5)\times 10^{-11}$ eV, or by a $100$-m instrument if $\mu\in(3-6)\times 10^{-11}$ eV.
@@ I Introduction
This scenario can be realized with physics beyond-the-Standard-Model (BSM). For example, a large number of ultralight fields may occur as a result of the compactification of extra dimensions [17] [18]. One of these may be the QCD axion – the pseudoscalar boson proposed to solve the strong-CP problem [19][20][21]. The axion is a Goldstone boson of a spontaneously-broken global symmetry which acquires a small mass through non-perturbative effects. Its mass, $\mu$, is determined by the energy scale $f_{a}$ associated with the broken symmetry [18],
@@
The $1-4\ M_{\odot}$ range of BH masses is gradually being populated by microlensing candidates [27] and GW events such as GW190814 [7] and GW230529 [8]. Since the mass distribution for these objects is still unknown, we will limit our attention to stellar-origin BH’s with masses between $5$ and $20\ M_{\odot}$, typical of BH’s in X-ray binary systems.
@@ II Superradiant bound-states
which is the radial Schrödinger equation for a non-relativistic particle in a Coulomb potential $V(r)=\alpha/r$ – hence the moniker ‘gravitational atom’. The $\alpha^{2}/(2n^{2})$ term in Eq. II is precisely the ‘hydrogen atom’ solution to Eq. 10. The complete small-$\alpha$ solutions to the full Klein-Gordon equation have been computed order-by-order using the method of matched asymptotic expansions [16] (and we have provided an example of Eq. II in Figs. 1 & 2),
@@ III Gravitational waves from axion clouds
At a particle-physics level, GW production by axion clouds can be understood in terms of two processes: annihilation of two bosons to a single graviton (with the BH absorbing the recoil momentum), and downward-transitions between bound-states [33] [34]. However, just as superradiance is a purely classical kinematic effect, the GW emission can also be understood classically in terms of the cloud’s time-dependent quadrupole moment. That being said, the GW signals considered in this work correspond to the annihilation channel.
@@
where $F_{+}(t)$ and $F_{\crossproduct}(t)$ are the detector’s angular pattern functions, and the phase $\Phi(t)$ encodes the frequency evolution in the detector frame,
[…]
where $\dot{f}$ is the frequency derivative, quantities with a zero-subscript are their values at a reference time $T_{0}$, and $T(t)$ is the time at the Solar System barycenter, related to the time at the detector $t$ by the addition of the Rømer, Shapiro, and Einstein delays.
The amplitudes $a_{+/\crossproduct}$ are expanded in terms of spheroidal harmonics with spin-weight $s=-2$,
@@
Comparing Eqs. 22 and 23, we see that $\mathcal{A}_{\tilde{l}\tilde{m}}\propto\sqrt{d\tilde{E}/dt}$, allowing us to express $h_{0}^{(\tilde{l})}$ directly in terms of $d\tilde{E}/dt$. Restricting ourselves to the dominant mode $\tilde{m}=\tilde{l}=2l$, we obtain a closed-form solution for the characteristic amplitude, which we use without abandon to compute the GW amplitudes of the axion clouds resulting from our simulations (we will drop the superscript $(2l)$ henceforth),
@@ V Simulation procedure
Those BH’s with an extant cloud are assigned a location in the Milky Way (Eqs. 31 and 32). Earth is assigned to an arbitrary, but fixed, point on the circle of radius $8.3$ kpc in the Galactic midplane. For a cloud located at distance $d$, we check the inequality $ct_{r}>d$ to determine if there has been enough time for GW’s to propagate to Earth since the cloud formed. Those clouds for which $d>ct_{r}$ are presently unobservable, and we retain only those clouds for which $ct_{r}>d$. We summarize this section with the following flowchart:
@@ VI.1 Cloud populations
In all cases, the distribution of GW frequencies occurs below the nominal value $f_{0}\propto 2\mu$ due to the positive scaling of gravitational redshift with BH mass. This interpretation is confirmed by a plot of the source-frame GW frequencies $f_{\text{GW,s}}$ and initial BH masses $M_{i}$ for all extant clouds in a given simulation (Fig. 7, with $\mu=3\crossproduct 10^{-11}$ eV). For each scalar bound-state, there is a tight relationship, with more massive BH’s producing lower-frequency clouds.
@@
Figure 7: The source-frame GW frequencies $f_{\text{GW,s}}$ and initial BH masses $M_{i}$ for the extant clouds in the $3\times 10^{-11}$ eV simulation. The spread of frequencies below $f_{0}\propto 2\mu$ is a result of gravitational redshift.
@@ VI.2 Resolvable signals
where $\sqrt{S_{n}(f)}$ is the one-sided amplitude spectral density (ASD) of the detector noise (the ‘sensitivity curve’) evaluated at the GW frequency, and the trapping frequency of the levitated sensor is constant during the entire observation time. The LSD is an Earth-bound detector for which the observed signal, Eq. 18, experiences both amplitude modulation via the angular-dependence of the sensitivity, and phase modulation via the Earth’s daily (diurnal) rotation and orbital motion. These corrections require knowledge of both the Earth’s position and the source position to high accuracy. In an all-sky blind search for axion clouds, the source position is not known ahead of time, so a realistic search will require a large number of templates corresponding to many sky locations. A coherent search for $\mathcal{O}(10)$ kHz GW’s over the full observation time (‘fully-coherent search’) of four months is not computationally-feasible because of the need to take fine steps in parameter space. A ‘semi-coherent’ search, in which the observation time is divided into several segments, sacrifices sensitivity for a great reduction in the number of templates needed to perform a coherent search on each segment [51].
As a means of setting upper limits on the expected number of resolvable signals, we compute the SNR for the idealized case of a detector freely orbiting the Milky Way at the same radius as the Solar System (i.e. not attached to a planet or star system). There is a positive frequency derivative due to the gradual dissipation of the cloud which, however, is estimated to be too small to be detected [30]. The signal remains perfectly coherent over the full observation time, and a hypothetical search performed at the Solar System barycenter would simply involve Fourier-transforming the data and looking for lines in the power spectrum.
@@
The ‘loudness’ of a signal is determined primarily by the source distance. The distance, in turn, is a random variable determined by the randomly-assigned position vector (Eqs. 31 and 32) of the source. Thus, for a given set of extant clouds, the number of individually-resolved sources $N_{\text{res}}$ will vary each time we re-assign their position vectors. We estimate the mean & standard deviation of $N_{\text{res}}$ for a given population of extant clouds by laying them down in the Galaxy $N_{\text{reshuffle}}=100$ times and counting how many are resolvable in each ‘re-shuffling’. The mean & standard deviation are then computed as
@@ VI.3 Unresolved signals
A preliminary method for estimating the LSD’s sensitivity to the confusion foreground is to treat each pair $(f_{c},\ h_{\text{eff}})$ as if they were the frequency and amplitude of a hypothetical monochromatic signal whose corresponding effective SNR $\rho_{\text{eff}}$, computed via Eq. 34, is then compared to a threshold $\rho_{\text{t}}$. We continue to require $\rho_{\text{t}}=10$. The numerator and denominator of Eq. 34 $\big{(}h_{\text{eff}}\sqrt{T_{\text{coh}}}\ \text{and}\ \sqrt{S_{n}(f_{c})},\ \text{respectively}\big{)}$ are shown separately in Fig. 11, and their ratio (the SNR) is shown in Figs. 12, 13, and 14 for the $1-$, $10-$, and $100-$m instruments, respectively.
@@ VII Conclusion
Finally, we note the following limitations of this work, as well as directions for future work: First, since isolated BH’s have no EM counterpart, we do not know, ahead of time, the direction to these GW sources. Targeted & directed searches for axion clouds will, therefore, not be possible for isolated BH’s, and we must resort to blind all-sky searches. Doppler modulations of the GW frequency can be accounted for by including the source’s right ascension & declination as additional parameters in the signal model for Bayesian parameter estimation. To avoid the number of demodulation templates becoming prohibitively large, – we do not want the time required for data analysis to be greater than the four-month observation period – we will resort to semi-coherent techniques for constructing a detection statistic. A final complication might be that a source has a transverse proper motion large enough to change its sky location within the observation time.
In this work, we have used the SNR as a baseline detection statistic. Since we have not yet developed the full semi-coherent search pipeline, we do not yet know what is the most appropriate detection statistic for continuous monochromatic signals. Another outstanding question pertains to our scan strategy: Given that we will take data at each trapping frequency for 4 months, how finely should the set of frequencies be discretized? At what frequency should we begin our searches? Figs. 8 and 11 both indicate the $14-20$ kHz range to be the most optimistic, in terms of the expected number of resolvable signals and the strength of the confusion foreground, respectively.
Additionally, our treatment of the confusion foreground has not accounted for the intrinsic anisotropy of the signal: The axion clouds will be distributed throughout the disks and bulge of the Milky Way, so the strength of the foreground will vary over the sky in a complicated way. Searches for stochastic signals typically involve an ‘excess-power’ method, as well as cross-correlation between multiple detectors. Plans to build a second $1$-m instrument at UC Davis (in addition to the Northwestern detector) are in development, so while a single $1$-m detector might not have the requisite sensitivity, the prospects for a two- or multi-detector scheme are an exciting avenue of future study.
@@ Appendix A Superradiant bound-states
Strictly speaking, the radial and angular eigenvalues, $\xi$ and $\Lambda$, must be found simultaneously. Leaver’s method can also be applied to 48 [54], resulting in a continued-fraction condition analogous to 65. We then have two equations for the two unknowns.
Conveniently, we can reduce the root-finding problem to merely solving 65 by using the $Mathematica$ function ‘SpheroidalEigenvalue’. With the change of variable $z=\cos\theta$, and in terms of the following quantities,
@@
For our axion cloud simulations, we have needed to compute $\xi$ for bound-states up to, and including, $n=9$. As an example, we have plotted the real & imaginary parts of the $n=8$ bound-state in Figs. 15, 16 and 17; Figs. 15 and 16 are analogous to Figs. 1 and 2.